The following is a report from a BLS survey taker: There were 110 people in the houses I visited. 28 of them were children under the age of 16, 25 people had full-time jobs, and 14 had part-time jobs. There were 10 retirees, 5 full-time homemakers, 13 full-time students over age 16, and 3 people who were disabled and cannot work. The remaining people did not have jobs, but all said they would like one. Nine of these people had not looked actively for work for three months, however. Find the labor force, the unemployment rate, and the participation rate implied by the survey takers report. Labor force: 42 people Instructions: Enter your responses rounded to two decimal places. Unemployment rate: 7140% Participation rate: %
Answer #1

Number of employed workers = 25+14 = 39

Number of unemployed workers = 110-(39+28+10+5+13+3) = 12-9 = 3

Labor force = employed +unemployed = 39+3 =42

Unemployment rate = 3/42*100 = 7.14%

Participation rate = Labor force/Adult population*100 = 42/(110-28) = 42/82*100 = 51.22%